Johnson Company calculates its allowance for uncollectible accounts as 10% of its ending balance in gross accounts receivable. The allowance for uncollectible accounts had a credit balance of $30,000 at the beginning of 2021. No previously written-off accounts receivable were reinstated during 2021. At 12/31/2021, gross accounts receivable totaled $500,000, and prior to recording the adjusting entry to recognize bad debts expense for 2021, the allowance for uncollectible accounts had a debit balance of 55,000.

Required:
1.
What was the balance in gross accounts receivable as of 12/31/2020?
2. What journal entry should Johnson record to recognize bad debt expense for 2021?
3. Assume Johnson made no other adjustment of the allowance for uncollectible accounts during 2021. Determine the amount of accounts receivable written off during 2021.
4. If Johnson instead used the direct write-off method, what would bad debt expense be for 2021?
Answer #1

Solution 1:

Balance in gross accounts receivable as of 12/31/2020 = balance in allowance for uncollectible accounts / 10% = $30000/ 10% = $300,000

Soltion 2:

Date Account title Debit Credit
31-Dec-2021 Bad debt expense (500,000*10% + 55000) 105,000
105,000

Solution 3:

amount of accounts receivable written off during 2021 = Beginning balance in allowance account - ending unadjusted balance in allowance account = $30000 - (-$55000) = $85,000

Solution 4:

bad debt expense be for 2021 under direct write off = $85,000